Mathematics 244 – Linear Algebra Definitions Quiz 4 – February 23, 2007

Your Name:

A) (5) Define: The set B is a basis for the subspace W ⊂ R^{n}.

B) (5) True or False, and explain (briefly!): If A is a 3 × 6 matrix, then we could have rank(A) = 2 and dim Nul(A) = 2.